OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. (-5.5,6.1),(-2.5,3.1) a.1 b.-3/8 c.-1 d.3/8

OpenStudy (anonymous):

use the equation : slope=( y2-y1)/(x2-x1)

OpenStudy (anonymous):

whats the answer ill give you a medal

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

slope=(3.1-6.1)/(-2.5+5.5)=-1